mathslover (mathslover):

Indefinite Integral Find the value of \[\large{ I = \int \cfrac{1}{x(x^n + 1)} dx }\]

mathslover (mathslover):

@vishweshshrimali5 This is what I have tried yet : \(\cfrac{1}{x(x^n +1) } = \cfrac{A}{x} + \cfrac{B}{x^n +1}\) \(\cfrac{A(x^n) + A + Bx}{x(x^n + 1)}\) Comparing this with the numerator of \(\cfrac{1}{x(x^n+1)}\) , we get : A = 0 and B = 1

mathslover (mathslover):

So, it can be written as : \(\int \left( \cfrac{1}{x(x^n+1)} \right) = \int \left( \cfrac{1}{x^n + 1} \right) \)

OpenStudy (vishweshshrimali5):

Nope since, n is not given here you can not use such partial fractions.

mathslover (mathslover):

Oh.. My bad. So, that all above is wrong? What to do now?

OpenStudy (vishweshshrimali5):

One way is to use integration by parts but I will not prefer it.

OpenStudy (vishweshshrimali5):

I am going to put x as e^u.

OpenStudy (vishweshshrimali5):

Then I get, \[\large{\int\cfrac{e^u * du}{e^u(e^{nu} + 1)}}\]

OpenStudy (vishweshshrimali5):

OKay I think I got it

OpenStudy (vishweshshrimali5):

Just forget all this.

OpenStudy (vishweshshrimali5):

Multiply and divide by x^{n-1}

OpenStudy (anonymous):

\[I=\int\limits \frac{ x ^{n-1}dx }{ x^n \left( x^n+1 \right) }\] put \[x^n=t,nx ^{n-1}dx=dt,\]

mathslover (mathslover):

Okay, so, that becomes : \[I = \int \frac{\frac{dt}{n}}{t(t+1)}\]

mathslover (mathslover):

\[ I = \int \cfrac{dt}{n \times t(t+1)}\] @Zarkon can you guide how to go on from here?

mathslover (mathslover):

Should I use partial fractions now?

mathslover (mathslover):

\[ I = \cfrac{1}{n} \int \cfrac{dt}{t(t+1)}\]

mathslover (mathslover):

Okay. Got it, Thanks @vishweshshrimali5 and @surjithayer :-) Using partial fraction will be the best I think.

OpenStudy (anonymous):

you are correct partial fractions is the best option.